OpenStudy (anonymous):

Rodríguez plans to order from 20 to 26 books for her class. The prices of the books she plans to order range from $4.95 to $12.95 each. If the publisher charges a shipping fee of $0.50 per book ordered, which of the following is not a reasonable price for the books, including shipping? A$145 B$245 C$345 D$445

OpenStudy (jackellyn):

Since the maximum amount of books is 26 and the maximum amount charges is $12.95+$0.50 shipping then =26*(12.95+0.50) =26(13.45) =$349.70 Any amount over that would be unreasonable.